# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ Appendix B GraSP General Overview
We present Figure 2 to facilitate the interpretation of phase, step, and action categories throughout their corresponding IDs, used in figures and tables along the main body of this work. In addition, the dendrogram aims to show the correspondence between phases and steps. These relationships reflect the usual proceedings inside operating rooms, where every surgical procedure has a “textbook” division into phases, which have themselves a “textbook” division into steps. They are also the results of the surgical team’s efforts to condense their experience into defining a set of processes that span a certain amount of time. Finally, it is essential to highlight the correspondence between the idle phase and idle step, as it shows that GraSP respects the logical relations between phases and steps. The consistency in the order of appearances of the phases and steps across all videos demonstrates that the GraSP dataset accurately represents the surgical scene.
@@ C.2 Surgical Transitions
Figures 7 and 9 show the transitions between phases and steps. These figures show how phase 0 and step 0 are recurrent throughout all surgeries and represent the transition between one step and another and one phase and another. Lastly, both tables and figures exhibit similarity in the order and duration of each phase and step during the temporal progression of the surgeries, hence showing the intrinsic procedural and ordered nature of surgical procedures and proving the consistency in annotation processes for both steps and phase tasks.
The transition matrices in Figures 5 and 6 present the transition probabilities of phases and steps across the GraSP dataset. Each value represents the probability of going from the previous step/phase (rows) to the next step/phase (columns). They demonstrate the logical connections between the phases and the steps. For instance, Figure 5 shows that Developing the space of Retzius always follows the Right pelvic isolated lymphadenectomy, as defined by the surgical team. There can be a pause between the two phases, corresponding to an Idle phase. The Figures show that most phases and steps transition from or to an Idle and that steps present more variability in their transition probabilities as step categories can show repeatedly throughout the procedure. Additionally, Figures 8 and 10 show the complete boxplots of the distribution of the duration of each phase and step categories. Besides the previous analysis in the Dataset Statistics Section, we observe that phase categories present more outliers, tending to have more extensive durations, once again proving the temporal granularity difference between phases and steps, where phases tend to represent broader temporal segments.
@@ C.3 Co-Occurence Matrices
Figures 12 and 13 display the inter-task co-occurrence matrices for instruments-phases and instruments-steps. These matrices demonstrate the varying frequency of distinct instruments during different stages of the surgery. Intuitively, the Large Needle Driver (LND) has the highest co-occurrence with the reconstructive phases of the surgeries like the Ligation of the dorsal venous complex and Bladder neck reconstruction phases, which involve the use of the LND to maneuver the surgical needles and threats to suture, ligate and reconstruct tissue. Hence, the LND is also the most frequent in suturing steps like Pass suture to urethra or bladder neck and pull or tie suture. Similarly, the Bipolar Forceps (BF) and Monopolar Curved Scissors are present in most phases and steps but especially towards phases like Left or Right pelvic isolated lymphadenectomy, Bladder neck identification and transection, and in steps like Identification and dissection of the Iliac vein and artery and Seminal vesicle dissection, which are mostly related to tissue manipulation and cutting. Moreover, the Clip Applier (CA) has a very high co-occurrence with the Prostatic pedicle control phase and the Vascular pedicle control step, which require tissue clipping for bleeding control. Also, the Laparoscopic Graspers/Instruments (LG or LI) show mostly in phases and steps requiring the holding of tissue or retrieval bags. Thus, these matrices again demonstrate the strong correlations between different semantic levels of surgical scene understanding.
@@ Appendix D TAPIS performance in GraSP
Figures 15, 16, and 17 depict qualitative segmentation results from TAPIS on the GraSP dataset. As mentioned previously, TAPIS can perform accurate segmentation of complex images. In addition, Figure 15 shows that TAPIS can segment instances of varying sizes, including bigger instrument arms and small tips. We show examples of inaccurate predictions in Figure 17. We conduct an error mode analysis and observe cases where two instrument instances are predicted as one, given a high degree of overlap between them (see first row, upper left corner). Second, we observe cases where an instrument is classified as another, given a problematic view of the instrument. For example, the fifth row confuses Bipolar Forceps with Prograsp Forceps. These instruments are similar and difficult to distinguish from a profile view. Third, we observe cases where it wrongly identifies other objects as instruments. For instance, the third row identifies an instrument for recollecting waste as a Laparoscopic Instrument. Additionally, Tables 7 and 8 show the per-class results on instrument segmentation explained in the Instrument Segmentation Experiments Section.
@@
On the one hand, Figures 18 and 19 present qualitative outcomes of TAPIS in the phase recognition task. Notably, our model encounters challenges in recognizing phases with shorter durations. In instances where recognition occurs, the model tends to identify longer time segments than their actual duration. Conversely, phases with longer durations exhibit more accurate recognition by the model.
On the other hand, Figures 20 and 21 illustrate results in the steps recognition task. Notably, the model demonstrates noise in recognizing steps with longer durations while performing well in predicting short steps. Additionally, timestamps may experience shifts, indicating challenges in defining transitions between steps. Tables 9 and 10 present the per-class quantitative results on phases and steps recognition as explained in the Phases and Steps Experiments Section.
@@
Method $mAP$ (%) Phase categories (AP %)
0 1 2 3 4 5 6 7 8 9 10
SlowFast 65.90 ±$\pm 4.99$ 64.94 70.93 73.74 81.14 72.89 84.36 62.24 33.03 43.82 66.05 71.74
TAPIS-VST 64.72 ±$\pm 1.64$ 61.41 69.92 76.06 78.17 72.40 83.78 66.75 26.83 40.45 61.10 75.04
TAPIS 72.90 ±$\pm 1.65$ 70.61 78.57 77.99 87.07 82.30 86.70 72.00 27.99 57.79 79.85 81.09
Method $mAP$ (%) Step categories (AP%)
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
SlowFast 42.60 ±$\pm 3.57$ 70.40 64.62 34.93 34.61 67.62 33.09 38.24 31.51 54.98 26.05 64.07 0.93 49.35 22.32 37.11 21.02 76.26 17.08 50.18 76.56 23.61
TAPIS-VST 41.02 ±$\pm 1.63$ 66.01 65.29 31.83 34.64 76.90 33.82 48.87 31.78 69.09 18.54 67.03 2.99 50.76 12.82 32.50 12.28 46.02 11.34 55.36 79.22 14.30
TAPIS 48.36 ±$\pm 1.14$ 73.14 74.15 42.22 34.92 80.18 39.30 41.60 35.83 65.32 24.48 63.62 1.35 77.41 35.49 36.68 17.88 77.34 23.13 63.56 83.41 24.47
@@
TAPIS Frame (SwinL) 34.04 ±$\pm 2.40$ 55.56 13.65 22.68 11.30 70.78 4.53 13.82 18.25 33.16 5.21 77.76 63.12 62.66 24.10

### v3

@@ Appendix B GraSP General Overview
We present Figure 2 to facilitate the interpretation of phase, step, and action categories throughout their corresponding IDs, used in figures and tables along the main body of this work. In addition, the dendrogram aims to show the correspondence between phases and steps. These relationships reflect the usual proceedings inside operating rooms, where every surgical procedure has a “textbook” division into phases, which have themselves a “textbook” division into steps. They are also the results of the surgical team’s efforts to condense their experience into defining a set of processes that span a certain amount of time. Finally, it is essential to highlight the correspondence between the idle phase and idle step, as it shows that GraSP respects the logical relations between phases and steps. The consistency in the order of appearances of the phases and steps across all videos demonstrates that the GraSP dataset accurately represents surgical scenes.
@@
Fig. 3: Short-Term Tasks Annotation Diagram. Our annotation process for instrument segmentation and atomic action detection (short-term tasks) was done per frame on keyframes sampled at 35s intervals. First, four surgical data science experts annotated all instrument instances for surgical atomic action recognition on each keyframe independently (top-part of the diagram). In parallel, surgical instruments were segmented by a team of 17 annotators trained and supervised by the surgical data science experts (middle part of the diagram). Third, both annotation types were aligned through bipartite matching, resulting in a single dataset with consistent instance segments and bounding boxes (bottom part of the diagram). We provide further details on the annotation process in Subsection 3.3 of the main manuscript.
@@ C.2 Surgical Transitions
Figures 9 and 11 show the transitions between phases and steps. These figures show how phase 0 and step 0 are recurrent throughout all surgeries and represent the transition between one step and another and one phase and another. Lastly, both tables and figures exhibit similarity in the order and duration of each phase and step during the temporal progression of the surgeries, hence showing the intrinsic procedural and ordered nature of surgical procedures and proving the consistency in annotation processes for both steps and phase tasks.
The transition matrices in Figures 7 and 8 present the transition probabilities of phases and steps across the GraSP dataset. Each value represents the probability of going from the previous step/phase (rows) to the next step/phase (columns). They demonstrate the logical connections between the phases and the steps. For instance, Figure 7 shows that Developing the space of Retzius always follows the Right pelvic isolated lymphadenectomy, as defined by the surgical team. There can be a pause between the two phases, corresponding to an Idle phase. The Figures show that most phases and steps transition from or to an Idle and that steps present more variability in their transition probabilities as step categories can show repeatedly throughout the procedure. Additionally, Figures 10 and 12 show the complete boxplots of the distribution of the duration of each phase and step categories. Besides the previous analysis in the Dataset Statistics Section, we observe that phase categories present more outliers, tending to have more extensive durations, once again proving the temporal granularity difference between phases and steps, where phases tend to represent broader temporal segments.
@@ C.3 Co-Occurence Matrices
Figures 14 and 15 display the inter-task co-occurrence matrices for instruments-phases and instruments-steps. These matrices demonstrate the varying frequency of distinct instruments during different stages of the surgery. Intuitively, the Large Needle Driver (LND) has the highest co-occurrence with the reconstructive phases of the surgeries like the Ligation of the dorsal venous complex and Bladder neck reconstruction phases, which involve the use of the LND to maneuver the surgical needles and threats to suture, ligate and reconstruct tissue. Hence, the LND is also the most frequent in suturing steps like Pass suture to urethra or bladder neck and pull or tie suture. Similarly, the Bipolar Forceps (BF) and Monopolar Curved Scissors are present in most phases and steps but especially towards phases like Left or Right pelvic isolated lymphadenectomy, Bladder neck identification and transection, and in steps like Identification and dissection of the Iliac vein and artery and Seminal vesicle dissection, which are mostly related to tissue manipulation and cutting. Moreover, the Clip Applier (CA) has a very high co-occurrence with the Prostatic pedicle control phase and the Vascular pedicle control step, which require tissue clipping for bleeding control. Also, the Laparoscopic Graspers/Instruments (LG or LI) show mostly in phases and steps requiring the holding of tissue or retrieval bags. Thus, these matrices again demonstrate the strong correlations between different semantic levels of surgical scene understanding.
@@
Fig. 17: TAPIS training diagram. We adopt a three-stage training process. First, we train our video feature extractor to address phase and step recognition jointly. We use a combined long-term loss ($\mathcal{L}_{longterm}$), consisting of adding the cross-entropy loss ($\mathcal{L}_{CE}$) of each task multiplied by a per task coefficient ($\lambda$) We observed that the values of $\lambda_{phases}$ and $\lambda_{steps}$ had a very significant impact on the performance of each task and could significantly enhance the performance of one task while strongly reduce the other. However, we found $\lambda_{phases}=1$ and $\lambda_{steps}=0.5$ to achieve the best balance. Second, we independently train Mask2Former (our instrument segmentation baseline) for instrument type instance segmentation on separate keyframes using its original segmentation loss ($\mathcal{L}_{Mask2Former}$). Third, we used the pretrained weights of stages 1 and 2 and we finetuned the entire TAPIS architecture for instrument reclassification or atomic action detection. We froze the segmentation baseline and stored its precalculated region proposals and features to avoid training overhead. Even though we observed that calculating region proposals on the fly during training yielded better performance, it significantly increased training FLOPS and speed, so we opted to employ precalculated regions.
@@ Appendix D TAPIS performance in GraSP
D.1 Dataset Splits for RARP45
Given the absence of a public test set in the original RARP-45 dataset, we create our splits, using 27 videos for training and 9 videos for evaluation. To ensure the integrity of the splits, we maintain a consistent class distribution with the original training set. Specifically, for the test split, we select videos 2, 3, 4, 6, 7, 9, 14, 17, 23, 28, 32, and 33, while we use the remaining videos for the training split.
@@ D.1 Dataset Splits for RARP45
Figures 18, 19, and 20 depict qualitative segmentation results from TAPIS on the GraSP dataset. As mentioned previously, TAPIS can perform accurate segmentation of complex images. In addition, Figure 18 shows that TAPIS can segment instances of varying sizes, including bigger instrument arms and small tips. We show examples of inaccurate predictions in Figure 20. We conduct an error mode analysis and observe cases where two instrument instances are predicted as one, given a high degree of overlap between them (see first row, upper left corner). Second, we observe cases where an instrument is classified as another, given a problematic view of the instrument. For example, the fifth row confuses Bipolar Forceps with Prograsp Forceps. These instruments are similar and difficult to distinguish from a profile view. Third, we observe cases where it wrongly identifies other objects as instruments. For instance, the third row identifies an instrument for recollecting waste as a Laparoscopic Instrument. Additionally, Tables 7 and 8 show the per-class results on instrument segmentation explained in the Instrument Segmentation Experiments Section.
@@
On the one hand, Figures 21 and 22 present qualitative outcomes of TAPIS in the phase recognition task. Notably, our model encounters challenges in recognizing phases with shorter durations. In instances where recognition occurs, the model tends to identify longer time segments than their actual duration. Conversely, phases with longer durations exhibit more accurate recognition by the model.
On the other hand, Figures 23 and 24 illustrate results in the steps recognition task. Notably, the model demonstrates noise in recognizing steps with longer durations while performing well in predicting short steps. Additionally, timestamps may experience shifts, indicating challenges in defining transitions between steps. Tables 10 and 11 present the per-class quantitative results on phases and steps recognition as explained in the Phases and Steps Experiments Section.
@@
TAPIS Frame (SwinL) 35.46 ±$\pm 0.34$ 60.26 13.93 22.43 13.77 71.64 5.31 22.02 17.60 31.46 5.69 76.80 65.97 64.27 27.51
Method mAP (%) 0 1 2 3 4 5 6 7 8 9 10
SlowFast $68.44\pm 0.10$ $72.89$ $67.83$ $66.63$ 80.53 $82.37$ $75.49$ $72.59$ 29.30 $56.29$ $71.78$ $75.02$
TAPIS-VST $70.60\pm 1.59$ $69.15$ 81.23 78.08 $76.49$ $79.45$ 80.84 73.73 $26.80$ $54.27$ $77.17$ 79.44
TAPIS $\textbf{71.36}\pm 1.33$ 73.16 $79.86$ $73.57$ 81.76 86.91 $78.84$ $71.88$ $24.76$ 56.66 80.42 $77.11$
Method mAP (%) 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20
SlowFast $43.86\pm 1.79$ $70.15$ $63.50$ $30.73$ $38.93$ $79.18$ $52.68$ $29.61$ $9.60$ $65.41$ $22.56$ $81.62$ $1.10$ $54.41$ $38.30$ $44.37$ $29.31$ $59.94$ $7.90$ $36.15$ $58.09$ $47.62$
TAPIS-VST $47.17\pm 2.65$ $70.34$ $57.01$ $31.36$ $31.64$ $78.92$ $55.98$ $45.20$ 11.39 74.97 $26.01$ 81.68 7.22 $79.52$ $36.12$ $45.29$ $33.58$ $49.52$ $8.30$ $38.56$ 79.23 $48.98$
TAPIS $\textbf{50.74}\pm 2.53$ 73.11 66.26 36.36 37.93 80.85 67.14 46.68 $10.11$ $71.81$ 27.10 $78.24$ $0.51$ 89.81 $\textbf{37.9}3$ 46.99 40.47 67.88 11.70 44.23 $76.47$ 53.35